Circular No. 6444

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) P.-O. Lagage and P. Galdemard, CEA, DSM/DAPNIA, and Service

d’Astrophysique, Centre d’Études de Saclay; and D. Bockelée-Morvan, Observatoire de Meudon, report: “Comet C/1995 O1 was observed on July 8.5 UT with the 3.6-m Canada-France-Hawaii Telescope (+ CAMIRAS, the Saclay 10-μm camera). Images (pixel resolution 0′′.3) revealed a northeast extension of the coma up to 35′′ from the nucleus (limited by the 60′′ × 40′′ field-of-view). The presence of a strong silicate feature with a peak at 11.2 μm, characteristic of crystalline olivine, is deduced from the following photometric measurements made over a square field of 24′′ × 24′′: 7.8–8.8 μm, 1.8 Jy; 9.4–9.75 μm, 6.4 Jy; 10.45–10.85 μm, 10.3 Jy; 11–11.45 μm, 17.9 Jy; 11.7–12.2 μm, 11.1 Jy; 11.8–13.1 μm, 13 Jy. The continuum color temperature is 185 K.”

X-RAY TRANSIENT NEAR SMC X-1 G. Clark, Massachusetts Institute of Technology; and J. Woo, Harvard-

Smithsonian Center for Astrophysics, report: “In ROSAT PSPC observations of SMC X-1 during 1991 Oct. 7–8, we have discovered a transient binary x-ray pulsar with a frequency of 2.41 Hz and strong quasiperiodic oscillations centered near 0.4 Hz. It lies so close to SMC X-1 that it causes no discernible distortion of the image. The power density of the combined flux at 2.41 Hz is approximately equal to that at 1.41 Hz and 2.82 Hz due to SMC X-1. Epoch folding shows a systematic decrease in the pulse frequency corresponding to a change in the radial component of velocity of ∼ 200 km/s during 1.0 day. No evidence of the transient is present in ROSAT observations of SMC X-1 on 1992 Oct. 1–2 and 1993 June 3–4 and in Ginga observations on 1989 July 29–Aug. 3. The presence of a strong transient so near SMC X-1 evidently requires caution to avoid confusion in the analysis of observations aimed at SMC X-1. Stars near Sanduleak 160 should be observed for evidence of binarity with a period in the range 2–4 days.”

SUPERNOVA 1996an IN NGC 1084 P. Garnavich, A. Riess, and R. Kirshner, Center for Astrophysics, re-

port that P. Berlind obtained CCD images of NGC 1084 on July 30.45 UT, using the F. L. Whipple Observatory’s 1.2-m telescope, that yield the following magnitudes: B ∼ 16.69, V ∼ 15.51, R ∼ 14.97, I ∼ 14.61.

1996 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green